# FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pre-Training

### v1

@@ 5 Related Works
When pre-training language models, a large amount of text corpus is often crawled from the internet. However, several studies Li et al. (2023); Zhou et al. (2024) suggest that high-quality data is more beneficial to the model’s performance. To select high-quality data, a common strategy involves utilizing rules crafted by experts Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and removing duplicate sentences Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). However, they often fall short in effectively selecting high-quality data based on semantic content. An alternative approach involves utilizing a target data source or proxy model Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024).
Training a classifier is a more straightforward method Du et al. (2022); Gururangan et al. (2022); Zhang et al. (2024); Wettig et al. (2024); Sachdeva et al. (2024). Du et al. (2022) implemented a logistic regression binary classifier to score the data, while some studies train more complex scorers Zhang et al. (2024); Sachdeva et al. (2024). Additionally, QuRating Wettig et al. (2024) trains multiple raters with a finer-grained approach to analyze the contribution of data to model performance improvement from different dimensions. The methods above primarily focus on selecting data from a single aspect. While QuRating has developed raters across multiple dimensions, it has not thoroughly explored how to integrate these raters, which is meticulously addressed in our research.
@@ 6 Conclusion
We develop a method that flexibly integrates multiple raters to assess the comprehensive quality of data for effective pre-training of LLMs. First, we align the ratings from different dimensions, thereby mapping the ratings to a unified score space. Next, we introduce orthogonality to adjust the weights of different raters. Due to the variation in orthogonality across different rating ranks, we further propose a progressive approach to perform fine-grained data selection. Experimental results on multiple benchmarks demonstrate that our integration method achieves outstanding performance, with a faster convergence rate compared to baselines.
@@ 7 Limitations and Future Works
Verification on larger models and more massive datasets
We’ve conducted extensive experiments on models of 1.3B and 3B in size and datasets ranging from 20 to 200B tokens. However, the applicability and effectiveness of our method on larger-scale models (such as models with more than 7B parameters) and larger, more diverse datasets have not been fully verified. Future studies should broaden our testing scope to fully evaluate our method’s scalability and versatility.
Appendix A Social Impact
@@ C.1 Analysis of the necessity of rating alignment
Multiple raters may employ different scales and criteria for assessing data quality, which can cause substantial problems if their ratings are integrated without appropriate standardization. For example, some raters may prioritize grammatical accuracy using a numerical scale, while others might assess semantic relevance using a percentage scale. Moreover, the rating thresholds distinguishing data that positively contribute to pre-training can significantly differ among raters. Without rating alignment, the integrated ratings can be misleading, inaccurately reflecting the actual quality of the data point. The subsequent examples and analyses underscore the importance of rating alignment for a reasonable data quality evaluation:
@@
  • Different Quality Thresholds. Even with ratings standardized to a common scale, we face the problem of varying quality thresholds distinguishing data that positively contribute to pre-training. For example, Rater A may deem ratings above 5 as high-quality, whereas Rater B may view ratings above 80% as high-quality. Suppose we standardize both ratings to a 0-1 scale, turning an 8 from Rater A into 0.8 and 85% from Rater B into 0.85. Despite this standardization, direct comparison of the two raters’ scores remains impractical due to their differing threshold values for differentiating data quality.
@@ C.2 Prompt for GPT-4o to compare data quality
Prompt for GPT-4o to compare data quality Compare two text excerpts and choose the text which contain more informative signal for pre-training a large-language model? An informative datapoint should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content. Aspects that should NOT influence your judgement: 1. The length of the text 2. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. The texts might have similar quality, but you should still make a relative judgement and choose the label of the preferred text. [Option A] {text a} [Option B] {text b} Now you have to choose between either A or B. Respond only with a single word.
@@ C.4 Ratings distribution illustration
The most intuitive approach to determine the orthogonality between two raters is to use the correlation coefficient between their rating distributions on a pre-training dataset. The orthogonality-correlation function needs to satisfy two key conditions:
@@
where $r(i,j)$ represents the correlation measure between raters $i$ and $j$. Specifically, it is quantified using the Pearson correlation coefficient between the score distributions of the two raters over a common dataset. In both Gaussian-based functions, the constant $c=\sqrt{\frac{1}{2\ln 2}}$ is determined by the boundary conditions. Figure 8 illustrates the 3 forms of orthogonality functions with respect to the correlation coefficient.
It’s worth noting that the experiments in D.2 demonstrate that the symmetrically processed gaussian function exhibits the best performance. Therefore, in the main experiments of this paper, we use the symmetrically processed gaussian function, i.e., Equation (14). Additionally, when integrating the ratings from two raters, their weights based on orthogonality are the same. Therefore, we only use the intrinsic reliability of the raters to weigh their ratings.
@@ D.2 Integration Baselines
D.3 Multi-dimension analysis prompt
We instruct the GPT-4o to evaluate the data selected though the raters, and here we show the prompts. For dimension, we consider four individual dimensions, the same as in Wetting et al. Wettig et al. (2024); and a comprehensive dimension:
@@ D.3 Multi-dimension analysis prompt
  • Does this text contain an informative signal for pre-training a large-language model? An informative datapoint should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content.
Prompts for GPT-4o evaluation You are a data annotation expert. Here is a piece of text for you to judge: {condition} Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. Here is the text: [TEXT BEGIN] {text} [TEXT END] Please follow the question order to respond. For answer, only respond yes or no. Return the results for each question in the following json format:
@@ E.1 Results of different combinations
We present the results of different rater combinations in Table 5. The average scores of FIRE (W.S.+R.E.+F.T.) and FIRE (W.S.+R.E.+E.V.) are 49.7 and 49.4 respectively, both surpassing the results obtained from combining any two of their individual raters. Furthermore, FIRE (W.S.+R.E.+F.T.+E.V.) achieves even better performance.
@@ E.4 Effect of Sample/top-K
(1) Our findings reveal that direct top-K selection outperforms sampling, further affirming the efficacy of our integration method. Additionally, Wettig et al. Wettig et al. (2024) posits that sampling enhances data diversity, which is beneficial for model learning. Our results indicate that the top-K scores are higher than the sampling scores, demonstrating that the top-ranked data according to our integration rating exhibit a broader distribution rather than being concentrated in a single domain.
(2) Contrary to the optimal value of 2 suggested by Wettig et al. Wettig et al. (2024), our analysis indicates that a relatively smaller $\tau$ value yields optimal results. A smaller $\tau$ accentuates the impact of ratings, suggesting that our method effectively selects higher-quality data. Furthermore, our multi-dimensional approach also accounts for the diversity of data types, ensuring a more comprehensive evaluation.

### v2

@@ 5 Related Works
When pretraining language models, a large amount of text corpus is often crawled from the internet. However, several studies Li et al. (2023); Zhou et al. (2024) suggest that high-quality data is more beneficial to the model’s performance. To select high-quality data, a common strategy involves utilizing rules crafted by experts Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and removing duplicate sentences Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). However, they often fall short in effectively selecting high-quality data based on semantic content. An alternative approach involves utilizing a target data source or proxy model Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024).
Training a classifier is a more straightforward method Du et al. (2022); Gururangan et al. (2022); Zhang et al. (2024); Wettig et al. (2024); Sachdeva et al. (2024). Du et al. (2022) implemented a logistic regression binary classifier to score the data, while some studies train more complex scorers Zhang et al. (2024); Sachdeva et al. (2024). Additionally, QuRating Wettig et al. (2024) trains multiple raters with a finer-grained approach to analyze the contribution of data to model performance improvement from different dimensions.
The methods above primarily focus on selecting data from a single aspect. While QuRating has developed raters across multiple dimensions, it has not thoroughly explored how to integrate these raters, which is meticulously addressed in our research.
@@ 6 Conclusion
We propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, enabling a comprehensive assessment of data quality across various dimensions. Initially, we align the ratings from different dimensions, mapping them to a unified score space. We then introduce orthogonality to adjust the weights of different raters. Given the variation in orthogonality across different rating ranks, we propose a progressive approach for fine-grained data selection. Experiments on the SlimPajama dataset reveal that FIRE outperforms other data selection methods, significantly enhancing the pretrained model’s performance across a wide range of downstream tasks. It achieves a 2.9% average performance improvement over Random and reduces the FLOPs required to reach a certain performance level by more than half.
@@ 7 Limitations and Future Works
Appendix A Ethical Considerations
@@ C.1 Analysis of the necessity of rating alignment
Multiple raters may employ different scales and criteria for assessing data quality, which can cause substantial problems if their ratings are integrated without appropriate standardization. For example, some raters may prioritize grammatical accuracy using a numerical scale, while others might assess semantic relevance using a percentage scale. Moreover, the rating thresholds distinguishing data that positively contribute to pretraining can significantly differ among raters. Without rating alignment, the integrated ratings can be misleading, inaccurately reflecting the actual quality of the data point. The subsequent examples and analyses underscore the importance of rating alignment for a reasonable data quality evaluation:
@@
  • Different Quality Thresholds. Even with ratings standardized to a common scale, we face the problem of varying quality thresholds distinguishing data that positively contribute to pretraining. For example, Rater A may deem ratings above 5 as high-quality, whereas Rater B may view ratings above 80% as high-quality. Suppose we standardize both ratings to a 0-1 scale, turning an 8 from Rater A into 0.8 and 85% from Rater B into 0.85. Despite this standardization, direct comparison of the two raters’ scores remains impractical due to their differing threshold values for differentiating data quality.
@@ C.2 Prompt for GPT-4o to compare data quality
Prompt for GPT-4o to compare data quality Compare two text excerpts and choose the text which contain more informative signal for pretraining a large-language model. An informative datapoint should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content. Aspects that should NOT influence your judgement: 1. The length of the text 2. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. The texts might have similar quality, but you should still make a relative judgement and choose the label of the preferred text. [Option A] {text a} [Option B] {text b} Now you have to choose between either A or B. Respond only with a single word.
@@ C.4 Ratings distribution illustration
The most intuitive approach to determine the orthogonality between two raters is to use the correlation coefficient between their rating distributions on a pretraining dataset. The orthogonality-correlation function needs to satisfy two key conditions:
@@
where $r(i,j)$ represents the correlation measure between raters $i$ and $j$. Specifically, it is quantified using the Pearson correlation coefficient between the score distributions of the two raters over a common dataset. In both Gaussian-based functions, the constant $c=\sqrt{\frac{1}{2\ln 2}}$ is determined by the boundary conditions. Figure 8 illustrates the 3 forms of orthogonality functions with respect to correlation coefficient.
It’s worth noting that the experiments in Appendix E.3 demonstrate that the symmetrically processed gaussian function exhibits the best performance. Therefore, in the main experiments of this paper, we use the symmetrically processed gaussian function, i.e., Equation (14). Additionally, when integrating the ratings from two raters, their weights based on orthogonality are the same. Therefore, we only use the intrinsic reliability of the raters to weight their ratings.
@@ D.2 Integration Baselines
D.3 Prompt for Multi-dimension Analysis
We instruct the GPT-4o to evaluate the data selected through the raters, and here we show the prompts. For dimension, we consider four individual dimensions, the same as QuRating Wettig et al. (2024); and a comprehensive dimension:
@@ D.3 Prompt for Multi-dimension Analysis
  • Does this text contain an informative signal for pretraining a large-language model? An informative data point should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content.
Prompts for GPT-4o evaluation You are a data annotation expert. You should judge that {condition} Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. Here is the text: [TEXT BEGIN] {text} [TEXT END] Please follow the question order to respond. For answer, only respond yes or no. Return the results for each question in the following json format:
@@ E.1 Results of different combinations
We present the results of different rater combinations in Table 5. The average scores of FIRE (W.S.+R.E.+F.T.) and FIRE (W.S.+R.E.+E.V.) both surpass the results obtained from combining any two of their individual raters. Furthermore, FIRE (W.S.+R.E.+F.T.+E.V.) achieves even better performance.
@@ E.4 Effect of Sample/top-K
(1) Our findings reveal that direct top-K selection outperforms sampling, further affirming the efficacy of our integration method. Additionally, Wettig et al. (2024) posits that sampling enhances data diversity, which is beneficial for model learning. Our results indicate that the top-K scores are higher than the sampling scores, demonstrating that the top-ranked data according to our integration rating exhibit a broader distribution rather than being concentrated in a single domain.
(2) Contrary to the optimal value of 2 suggested by Wettig et al. (2024), our analysis indicates that a relatively smaller $\tau$ value yields optimal results. A smaller $\tau$ accentuates the impact of ratings, suggesting that our method effectively selects higher-quality data. Furthermore, our multi-dimensional approach also accounts for the diversity of data types, ensuring a more comprehensive evaluation.